OpenStudy (anonymous):

\[\frac{(x ^{-5}x ^{4})^{4} }{ (x ^{4}x ^{-6})^{-5}}\]

OpenStudy (anonymous):

How do i do this...Yes it may be simple but for some reason i can't figure it out.

TheSmartOne (thesmartone):

Hint: First use this formula \(\sf(a^b)^c = a^{b\times c}\) Then use this one \(\sf\Large\frac{a^b}{a^c}=a^{b-c}\)

OpenStudy (anonymous):

I know that

TheSmartOne (thesmartone):

also these formulas: \(\sf a^b\times a^c = a^{b+c}\) \(\sf\Large {a^{-b}}=\frac{1}{a^b}\)

OpenStudy (anonymous):

yes i understand that...i just can't seem to get the correct answer to this...

OpenStudy (phi):

can you simplify \[ x ^{-5}x ^{4} \]? (add the exponents)

OpenStudy (anonymous):

x^-1

OpenStudy (phi):

yes now we have up top \[ (x^{-1})^4 \] now use the rule multiply exponents to simplify that

OpenStudy (anonymous):

X^-4

OpenStudy (phi):

now let's do the bottom. first \[ (x ^{4}x ^{-6}) \]

OpenStudy (anonymous):

x^-2

OpenStudy (phi):

now (x^-2)^-5

OpenStudy (anonymous):

x^10

OpenStudy (phi):

so we have \[ \frac{x^{-4}}{x^{10} }\]

OpenStudy (anonymous):

so 1/ x^14

OpenStudy (phi):

you can do that two ways: "flip" x^-4 and put it in the bottom but with x^4 then combine x^4*x^10 (in the bottom) or do -4 - 10 (subtract the exponents because we divide)

OpenStudy (phi):

yes 1/x^14 this can also be written x^(-14)

OpenStudy (anonymous):

ok thank you!